# Coherent Control of an Optical Quantum Dot Using Phonons and Photons

### v1

@@ III Resolving phonon-assisted excitation dynamics
IV Comparing pulsed and CW pumping schemes
Fig. 4 summarizes our experimental (open markers) and numerical (solid/dashed curves) results, where we plot maximal enhancements as a function of $g$. We theoretically consider three idealized and simple pulse shapes — square, “exponential”, and triangle, as illustrated in the top panels — and TLSs with both radiative relaxation only (solid curves) and additional pure dephasing (dashed curves); see Methods. “Exponential” and “square” correspond approximately to our two experimental cases, “exponential” referring to the exponentially decaying tail. With optical pulse lengths comparable to the TLS lifetime of $\approx$1 ns, triangular and “exponential” pulse shapes provide very similar performance, and both provide approximately ten-fold greater maximal enhancement than square pulses. The importance of both (1) a gradually decreasing pulse intensity and (2) lifetime-limited TLS dephasing are clear. Additional dephasing over timescales comparable to the QD lifetime — and thus the pulse duration — rapidly reduces maximal enhancements. Our experimental results evidence some form of dephasing in addition to radiative decay. Nonetheless, the experimentally obtained improvement between pulse shapes corresponds well to the ten-fold improvement seen in calculations. This suggests that our results are limited by slow spectral fluctuations rather than pure dephasing at the timescale of the QD’s lifetime. That is, our experimental occupancies are slightly “smeared” as the detuning effectively varies between successive detection events. This will be largely alleviated by investigating QDs with narrower measured linewidths.
[…]
@@ V Discussion
We experimentally demonstrated how coherent optical control methods available with QD optomechanical systems can be exploited to enhance phonon-assisted scattering in applications such as microwave-to-optical transduction. Specifically, time-correlated single-photon counting measurements reveal excitation dynamics in a semiconductor QD that depend on the presence of microwave-frequency phonons. Pulse durations exist where background scattering from direct excitation is greatly reduced, improving the relative rate from phonon-assisted scattering. This signal-to-background ratio depends on the basic features of the optical pulse, including amplitude, detuning, shape, and phase. We explored three very simple distinct pulse shapes and demonstrated that pulses with a gradually decreasing intensity are beneficial. More complicated pulse shapes involving frequency, amplitude and phase shaping may be developed to further improve the performance.
@@
More generally, this work shows that with the QD, new parameters are available for interacting with the mechanical system. For example, in addition to pump intensity, the performance now depends on details of the varying pump — including the amplitude and phase (not explored in this work) — at nanosecond timescales. In fact, with this time-domain control, there is no fundamental reason for matching the detuning to the mechanical mode frequency as is needed with bosonic resonators [2]. Using additional detuned pumps may offer another level of control, similar to previously demonstrated “swing-up” schemes [21, 32, 50]. Since the QD’s lifetime sets the timescale of state preparation and readout, average optical pump powers used for these protocols are at the nanowatt scale and this optical power is not stored in a resonator, reducing absorption-induced heating which is a major concern in optomechanical crystals [64, 13, 18].
Comparing our experimental results to numerical calculations suggests that our measured QD experiences slow spectral fluctuations that decrease the experimentally obtained enhancements. This is likely reflected in our measured QD linewidth of 800 MHz, but additional characterization would be needed to disentangle pure dephasing at the timescale of the QD’s lifetime from slow (sub-MHz) spectral wandering [60]. Modern optimized charge-controlled QDs regularly exhibit near-lifetime-limited linewidths [59, 57, 65], and we thus anticipate that performance comparable to our idealized numerical calculations is possible. Overall improved performance is anticipated from p-i-n charge-controlled QDs which offer greater charge and frequency stability, greater optical pumping efficiency, higher photon collection efficiency, and reduced pump background. Although our numerical calculations do not take into account the finite pump bandwidth, this is likely to play only a small role in experiments since the ideal pulse durations are longer than 1 ns, corresponding to transform-limited bandwidths of $\approx$350 MHz which are small compared to our detuning of $\approx$3.5 GHz. Larger issues likely arise from small polarization drifts in the experimental system which may change the field amplitude and temporal pulse shape that the QD experiences over the course of the 30-60 second integration times used in our experiments. Future efforts should involve optimizing QD performance in application-relevant mechanical resonators and exploring the broad parameter space of optical pumping schemes for improving the enhancements described here.
@@ VIII.1 System Hamiltonian
The Hamiltonian describing our system, in a frame rotating with the optical pump at frequency $\omega_{\text{pump}}$, is given by
@@
VIII.2 Experimental Setup
@@
All of our QD optical measurements were performed using a 5 K optical cryostat with an internal cryogenic 0.81 NA objective. The resonant pump beam is shaped outside the cryostat using single-mode optical fiber components and free-space optical components. The resonant pump beam was a tunable external-cavity diode laser with a 10 kHz linewidth. Pump beams illuminate the QD through the objective. QD luminescence was collected by the same objective and isolated from the pump predominantly by spectral filtering (for measurements involving a nonresonant pump) or polarization filtering (for measurements involving a resonant pump) with at least 106 pump rejection. Luminescence was collected into single-mode optical fiber and various frequency- and time-domain measurements were performed using procedures detailed below. The QD bias voltage was held constant for all measurements shown in the manuscript. The QD emission wavelength (and therefore the resonant pump wavelength) used in this work was 975.934 nm.
VIII.3 Continuous-wave Resonance Fluorescence Spectra
@@
Continuous-wave (CW) resonance fluorescence (RF) spectra (e.g. Fig. 2) were recorded by slowly sweeping the frequency of a tightly focused resonant pump beam over the QD’s exciton resonance while holding the QD bias constant. Luminescence counts at each frequency were recorded using a superconducting nanowire single photon detector (SNSPD) mounted in a dilution refrigerator. A pump power of approximately 2 nW at the sample was used, corresponding to approximately 1/10 the measured saturation level of the QD. For manuscript Fig. 2d (upper panel), such spectra were recorded without SAW driving and with SAW driving using an external coherent microwave source at approximately 3.518 GHz and external microwave driving powers between -50 dBm and -35 dBm. External microwave driving powers were converted to coupling rates $g$ by a series of calibration data where the modulation index $\chi$ was fit for each microwave driving power and $g$ was derived by the relationship $g$=$\omega_{\text{SAW}}\chi$. The calibration data are shown in the Supplementary Information.
@@
For manuscript Fig. 2d (lower panel), the “resonant” pump was held at a constant frequency detuning of $\Delta$=$-\omega_{\text{SAW}}$. A pump power of approximately 5 nW was used. The frequency spectrum of scattered photons was measured by passing the collection port through a tunable Fabry-Perot etalon filter with a 600 MHz linewidth before being counted on an SNSPD. In manuscript Fig. 2d (lower panel), such spectra were again measured without SAW driving and with SAW driving at 3.518 GHz and the same powers as in manuscript Fig. 2d (upper panel).
VIII.4 Pulsed Optical Measurements and Time-domain QD Dynamics
@@
The time evolution of the QD dynamics was measured using standard time-correlated single-photon counting (TCSPC) techniques with a two-channel picosecond event timer and TCSPC module. For “exponential” decaying pulses, optical pulses were generated by passing a CW “resonant” optical pump (same as described above) through two cascaded electro-optic modulators (EOMs). The EOMs were driven by amplified pulses from an approximately 100-ps square-wave electronic pulse generator (EPG). The EPG was externally triggered using an arbitrary waveform generator (AWG) at 38 MHz. The rise time, fall time and width of the output pulses were measured to be approximately 15 ps, 15 ps, and 130 ps respectively. The on/off contrast was measured to be better than $10^{4}$. The frequency bandwidth of the output pulses was measured to be 7 GHz. To decrease the bandwidth (and increase the pulse duration), the pulses were transmitted through a 600 MHz tunable Fabry-Perot etalon which was tuned for maximum transmission at the pulse’s center frequency. For square pulses, the EOMs were directly driven using a high-speed AWG (64 GS/sec) with approximately 30 ps rise time and pulses of arbitrary length. The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement). Measured pulse profiles shown in the manuscript were from the reflected signal and may contain system reflections that don’t illuminate the QD. That is, the pulse profile experienced by the QD may differ slightly from what we measured. This may partially account for the onset of the QD’s free evolution around 800 ps in Fig. 3a although the measured pulse continues beyond this time.
@@
The time distribution of the QD luminescence was measured by again detecting the luminescence with an SNSPD and time-correlating the counts on the TCSPC module with respect to a trigger signal from the AWG. We typically used 30-60 second total integration times, over which we collected typically several million total photons. This type of measurement was performed with a wide range of average optical input powers (between approximately 0.25 nW to 1500 nW). Pump powers were converted to resonant Rabi rates by looking at the resulting time-domain oscillations.
@@
Supplementary Information
@@ 1 System Hamiltonian
Our system comprises a two-level system (TLS) with frequency $\omega_{\text{QD}}$ driven at optical frequency $\omega_{\text{pump}}$ with a coupling (Rabi) rate $\Omega_{0}$. The TLS frequency is modulated at microwave frequency $\omega_{\text{SAW}}$ with a modulation depth $\chi$=$g/\omega_{\text{QD}}$. The Hamiltonian $\hat{H}$ is:

### v2

@@ III Experimentally resolving phonon-assisted excitation dynamics
IV Distinguishing Scattering Channels During Phonon-Assisted Pulsed Excitation
The spectrum of photons scattered from the coupled QD-SAW system carries information about the state of both the SAW resonator and the QD. In our time-resolved measurements presented above, we inferred the QD occupancy from the total number of photons scattered or emitted over all frequencies. Those measurements thus generally do not distinguish between various scattering mechanisms including, e.g., Rayleigh scattering and QD spontaneous emission. To address this, we prepare our system using two distinct red-detuned ($\Delta$=$-\omega_{\text{SAW}}$) square pulses (Fig. 4a). This state preparation is visualized through the QD dynamics both without SAW driving (gray curves) and with moderate SAW driving (orange curves). The scattered photons are then frequency resolved with filters of varying transmission bandwidths, $\delta f$ (Fig. 4b). Pulse durations are chosen to provide a maximum contrast between the two spectra; the shorter pulse (duration $\tau_{\text{pulse}}$$\approx$370 ps) ends at a maximum of the coherent zero-phonon Rabi cycle (Fig. 4a; left panel), and the longer pulse ($\tau_{\text{pulse}}$$\approx$500 ps) ends at the next minimum (right panel).
Spectra measured under the two different pulse scenarios with and without SAW driving are displayed in Fig. 4c. The first notable feature is the weak peak at the laser drive frequency ($\Delta$=$-\omega_{\text{SAW}}$). We attribute this to Rayleigh scattering, identified by the spectral widths which are characteristic of the optical pulses for each pulse type, and by the little variation from case to case. The resonance at $\Delta$=0 is more interesting. Comparing spectra with and without SAW driving, we identify components related to both the detuned zero-phonon excitation of the QD and the phonon-assisted excitation, depending on the preparation protocol. With the short pulse — which can be considered a “poor” choice for isolating optomechanical excitation — both spectra are almost completely dominated by the zero-phonon excitation process. With the longer pulse (right two panels) most of the photons are from the phonon-assisted excitation process (orange). These results are in fact the experimental realization of the process illustrated in Fig. 1c.
Figure 4: (a) Experimental time-resolved photon signal with red-detuned ($\Delta$=$-\omega_{\text{SAW}}$=$-2\pi\times$3.58 GHz) optical pulses of lengths $\tau_{\text{pulse}}$$\approx$370 ps (left panel) and $\tau_{\text{pulse}}$$\approx$500 ps (right panel). Gray curves: no SAW drive ($g$=0). Orange curves: SAW drive ($g$$\approx$1 GHz). (b) Schematic experimental setup for measuring the spectra shown in panels b) and c). After pulsed excitation with and without SAW driving, collected photons are passed through a tunable spectral filter with transmission bandwidth $\delta f$ and measured either by integrating over all photon arrival times (‘Int.’) or by recording arrival times with TCSPC. (c) Spectra measured with $\delta f$=25 MHz. Left two panels correspond to $\tau_{\text{pulse}}$=370 ps. Right two panels correspond to $\tau_{\text{pulse}}$=500 ps. Panels are split to highlight two important spectral features around $\Delta$=$-\omega_{\text{SAW}}$ and $\Delta$=0. The spectral width of each subpanel is identical (3.8 GHz). Data (dots) have been fit to Gaussian functions (solid curves). Text specifies the fit full-width at half maximum. (d) Time-resolved photon signal (same setup as in panel a) measured with $\delta f$=1 GHz around $\Delta/2\pi$=--3.58 GHz (red curve) and $\Delta$=0 (black curve). The gray region is the measured pump background. Insets show reconstructed spectra at the 0.5 ns (circle markers) and 2.0 ns (square markers).
More information is obtained by measuring the time distribution of collected photons after passing them through a $\delta f$=1 GHz spectral filter (Fig. 4d). This allows the pump and QD frequencies to be isolated while retaining temporal resolution on the order of $1/\delta f$$\approx$1 ns; time traces become “smeared” over about 1 ns. The decay of the $\Delta$=$-\omega_{\text{SAW}}$ component (red curve) follows the abrupt termination of the excitation pulse, while the $\Delta$=0 component (black curve) decays much more slowly, at approximately the QD lifetime. This further confirms Rayleigh scattering as the origin of most of the the red-detuned photons; QD preparation and subsequent luminescence accounts for the emission at $\Delta$=0. These different temporal dynamics are consistent with the spectral widths of the peaks measured in Fig. 4c. From this data, we make artificial reconstructions of the scattering spectrum at 0.5 ns (Fig. 4d; left inset, circle markers) and 2 ns (right inset; square markers). The spectrum varies in time during and after the pulsed excitation, largely resembling the QD at long times after the pump terminates, confirming the successful preparation of the QD state.
@@ V Discussion
We experimentally demonstrates and theoretically described phonon-assisted excitation dynamics in a semiconductor QD using detuned optical pulses and coherent phonon driving. Specifically, time-correlated single-photon counting measurements reveal excitation dynamics that depend on the presence or absence of microwave-frequency phonons. We showed that the QD can be prepared in an excited state preferentially by phonon-driven processes with strategic optical pulse shaping. That is, the specific dynamics are sensitive to the presence of phonons and details of the optical pulse, including amplitude, temporal envelope, detuning, and phase. Complementary frequency-domain measurements reveal time-dependent scattering channels related to Rayleigh scattering and luminescence, and verify the successful preparation of the QD excited state. These behaviors may be exploited for applications such as microwave-to-optical transduction where photon scattering unrelated to phonon-assisted processes is highly undesirable. In this context, we demonstrated that pulses with a gradually decreasing intensity are beneficial. More complicated pulse shapes involving frequency, amplitude and phase shaping may be developed to further improve this performance.
@@
More generally, this work shows that with the QD, new parameters are available for interacting with the mechanical system at nanosecond timescales. With this time-domain control, there is no fundamental reason for matching the detuning to the mechanical mode frequency as is needed with bosonic resonators [2]. Using additional detuned pumps may offer another level of control, similar to previously demonstrated “swing-up” schemes [22, 33, 55]. Since the QD’s lifetime sets the timescale of state preparation and readout, average optical pump powers used for these protocols are at the nanowatt scale and this optical power is not stored in a resonator, reducing absorption-induced heating which is a major concern in optomechanical crystals [63, 13, 17].
Comparing our experimental results to numerical calculations suggests that our measured QD experiences slow spectral fluctuations that hinder the ability to repeatedly prepare the QD in a desired state over many seconds. Modern optimized charge-controlled QDs regularly exhibit near-lifetime-limited linewidths [60, 58, 64], and we thus anticipate that performance comparable to our idealized numerical calculations is possible. Future efforts should involve optimizing QD performance in application-relevant mechanical resonators with micrometer-scale mode volumes and exploring the broad parameter space of optical pumping schemes for improving the enhancements described here.
@@ VIII.1 System Hamiltonian
The Hamiltonian describing our system, in a frame rotating with the optical pump at frequency $\omega_{\text{pump}}$, assuming a semiclassical optical and mechanical drive is given by
@@
VIII.2 Device Fabrication
Full device fabrication details are provided in the Supplementary Information.
VIII.3 Experimental Setup
@@
All of our QD optical measurements were performed using a 5 K optical cryostat with an internal cryogenic 0.81 NA objective. The resonant or detuned pump beam is shaped outside the cryostat using both single-mode optical fiber components and free-space optical components. The resonant pump beam was a tunable external-cavity diode laser with an approximate 10 kHz linewidth at 960 nm. Pump beams illuminate the QD through the objective. The QD signal was collected by the same objective and isolated from the pump predominantly by spectral filtering (for measurements involving a nonresonant or detuned pump) or polarization filtering (for measurements involving a resonant pump) with at least 107 pump rejection. Luminescence was collected into single-mode optical fiber and various frequency- and time-domain measurements were performed using procedures detailed below. The QD bias voltage was held constant for all measurements shown in the manuscript. The QD absorption/emission wavelength (and therefore the resonant pump wavelength) used in this work was 956.368 nm.
VIII.4 Continuous-wave Resonance Fluorescence Spectra
@@
Continuous-wave (CW) resonance fluorescence (RF) spectra (e.g. Fig. 2d) were recorded by slowly sweeping the frequency of a tightly focused resonant pump beam over the QD’s exciton resonance while holding the QD bias constant. Luminescence count rates at each frequency were recorded using a superconducting nanowire single photon detector (SNSPD) mounted in a dilution refrigerator. A pump power of approximately 2 nW at the sample was used, corresponding to approximately 1/10 the measured saturation level of the QD. For manuscript Fig. 2d (left panel), such spectra were recorded without SAW driving and with SAW driving using an external coherent microwave source at approximately 3.587 GHz and external microwave driving powers between -50 dBm and -35 dBm. External microwave driving powers were converted to coupling rates $g$ by a series of calibration data where the modulation index $\chi$ was fit for each microwave driving power and $g$ was derived by the relationship $g$=$\omega_{\text{SAW}}\chi$. The calibration data are shown in the Supplementary Information.
@@
For manuscript Fig. 2d (right panel), the pump was held at a constant frequency detuning of $\Delta$=$-\omega_{\text{SAW}}$. A pump power of approximately 5 nW was used. The frequency spectrum of scattered photons was measured by passing the collection port through a tunable Fabry-Perot etalon filter with a 600 MHz linewidth before being counted on an SNSPD. In manuscript Fig. 2d (right panel), such spectra were again measured without SAW driving and with SAW driving at 3.587 GHz and the same powers as in manuscript Fig. 2d (upper panel). Note that SAW driving frequencies varied by several MHz between measurements due to cavity frequency shifts over the duration of the measurements which spanned several days.
Further experimental details are provided in the Supplementary Information.
VIII.5 Pulsed Optical Measurements and Time-domain QD Dynamics
@@
The time evolution of the QD dynamics was measured using standard time-correlated single-photon counting (TCSPC) techniques with a two-channel picosecond event timer and TCSPC module. For “gradual decrease” pulses, optical pulses were generated by passing a CW detuned optical pump (same as described above) through two cascaded electro-optic modulators (EOMs). The EOMs were driven by amplified pulses from an approximately 100-ps square-wave electronic pulse generator (EPG). The EPG was externally triggered using an arbitrary waveform generator (AWG) at 34.085 MHz repetition rate. The rise time, fall time and width of the output pulses were measured to be approximately 15 ps, 15 ps, and 130 ps respectively. The on/off contrast was measured to be about $10^{5}$. The frequency bandwidth of the output pulses was measured to be 7 GHz. To decrease the bandwidth (and increase the pulse duration), the pulses were transmitted through a 600 MHz tunable Fabry-Perot etalon which was tuned for maximum transmission at the pulse’s center frequency. The resulting pulse shape is the “gradual decrease” shape shown in Fig. 3. For square pulses, the EOMs were directly driven using a high-speed AWG (64 GS/sec) with approximately 30 ps rise time and pulses of arbitrary length. The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement).
@@
The time distribution of the QD luminescence was measured by again detecting the luminescence with an SNSPD and time-correlating the counts on the TCSPC module with respect to a trigger signal from the AWG. We typically used 30-60 second total integration times, over which we collected typically several million total photons. This type of measurement was performed with a wide range of average optical input powers (between approximately 0.25 nW to 300 nW). Pump powers were converted to resonant Rabi rates by looking at the resulting time-domain oscillations with zero detuning.
@@
VIII.6 Spectral measurements with pulsed excitation
Spectral measurements shown in Fig. 4c,d were performed by collecting and passing the QD signal through two distinct frequency-tunable Fabry-Perot etalon filters. The pump configurations were identical in both cases, specified in the manuscript. For Fig. 4c, we used an etalon with a 25 MHz transmission bandwidth and present the data at each frequency (voltage) value. Fig. 4d, we used an etalon with a 1 GHz transmission bandwidth and present the resulting time-domain data for only two different transmission frequency windows corresponding to the (detuned) pump frequency and the QD frequency.
Supplementary Information:
Coherent Control of an Optical Quantum Dot Using Phonons and Photons
@@ 1 System Hamiltonian
1.1 General description
Our system comprises a two-level system (TLS) with frequency $\omega_{\text{QD}}$ driven at optical frequency $\omega_{\text{pump}}$ with an amplitude described by the coupling (Rabi) rate $\Omega_{0}$. The TLS frequency is modulated at microwave frequency $\omega_{\text{SAW}}$ with a modulation depth $\chi$=$g/\omega_{\text{QD}}$. Here, we treat the optical drive and the mechanical drive semiclassically, which is appropriate for moderate SAW driving powers as used in our experiments. In this case, the phonon field amplitude is contained in the rate $g$, which can be related to the single-phonon coupling rate $g_{0}$ by $g=g_{0}\sqrt{n}$ when $n$ phonons occupy the resonator. The Hamiltonian $\hat{H}$ is:
@@ 1.1 General description
When treating the phonon mode quantum mechanically, single-phonon scattering terms manifest with coupling rates $\Gamma_{-}=g_{0}\Omega_{0}\sqrt{n+1}/\omega_{\text{SAW}}$ and $\Gamma_{+}=g_{0}\Omega_{0}\sqrt{n}/\omega_{\text{SAW}}$, corresponding to removing and adding single phonons from the resonator, respectively. This originates from a perturbative expansion of the coupling Hamiltonian in the normalized rate $g/\omega_{m}$, and so applies formally only to small phonon occupancies. See reference 53 of the manuscript for further details. For large $n$, these rates are similar and we refer to them both simply as $\Gamma$ in the manuscript.
@@
1.2 Calculation details
In Figs. 1b,c of the manuscript we present the dynamics in two different pictures. In both panels, we assume an initial state $|g,n\rangle$, a total evolution time of 1.7 ns, and no dephasing. In Fig. 1b, we calculate the occupancy of the state $|e,n\rangle$ by calculating TLS dynamics between states $|g,n\rangle$ and $|e,n\rangle$ with Rabi rate $\Omega_{0}/2\pi$=1 GHz and $\Delta/2\pi$=$-\omega_{\text{SAW}}/2\pi$=--3.5 GHz (black curve). We then calculate the occupancy of the state $|e,n-1\rangle$ by calculating TLS dynamics between states $|g,n\rangle$ and $|e,n-1\rangle$ with the renormalized interaction rate $\Gamma$=$g_{0}\Omega_{0}\sqrt{n}/\omega_{\text{SAW}}$ with $\Delta$=0, $\Omega_{0}/2\pi$=1 GHz and $g_{0}\sqrt{n}$=$2\pi\times$1 GHz. These simplifications are largely for illustrative purposes, although they may be considered accurate in the artificial case of no dephasing. In Fig. 1c, we ignore the discrete phonon levels and calculate the QD’s excited state occupancy by calculating the dynamics between the bare states $|g\rangle$ and $|e\rangle$ with $\Omega_{0}/2\pi$=1 GHz, $\Delta/2\pi$=$-\omega_{\text{SAW}}/2\pi$=--3.5 GHz and optomechanical coupling rates $g$=0 (black trajectory) and $g/2\pi$=2 GHz (red trajectory). Calculations in the manuscript Fig. 3 are like the ones shown in Fig. 1c. We assume an initial QD state $|g\rangle$, detuning $\Delta$=$-\omega_{\text{SAW}}$$\approx$--3.5881 GHz, Rabi rates $\Omega_{0}$ corresponding to measured pump amplitudes (specified in the manuscript) and $g$ values corresponding to measured microwave powers (described further in the calibration procedure below).
1.3 Mechanical phase